# Concomitant Entanglement and Control Criticality Driven by Collective Measurements

### v1

@@ Acknowledgements.
We acknowledge helpful discussions with Wen-Wei Ho, Nishad Maskara, and Sarang Gopalakrishnan. We especially thank Sriram Ganeshan and Michael Buchhold for numerous inspiring conversations and collaboration on related topics. This work was supported in part by the National Science Foundation under grants DMR-2143635 (T.I.) and DMR-2238895 (J.H.W.). J.H.P. was supported by the Office of Naval Research grant No. N00014-23-1-2357. This work was initiated and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation Grant No. PHY-1607611. This work was also supported in part by the International Centre for Theoretical Sciences (ICTS) via participation in the program “Periodically and quasi-periodically driven complex systems” (code: ICTS/pdcs2023/6). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (J.H.P.).
Appendix A Control onto General Short Orbits
@@
Appendix B Finite-Size Decay of Observables
@@
To remove this decay, we add one more step to the chaotic circuit of Sec. II.1. After applying $U_{\rm PXP}$ and $U_{\rm sigma}(\theta)$, we apply a (conditional) bit flip gate of the form (9) on a single qubit, chosen at random each time the chaotic circuit is applied; The rationale for this is that a finite fraction of the wave function gets trapped on the vacuum orbit and if we introduce a perturbation that is sub-extensive and kicks us off of the orbit, the system will not fully decay in the putative volume-law phase (at small system sizes). The result is illustrated with dashed lines in Fig. 13, which display clear saturation as expected in a standard dynamical phase. While we have not verified that the transition persists in the presence of this modification, a similar strategy was employed for $\beta$-adic Rényi circuits in Refs. [7, 8, 9, 10], where a clear second-order phase transition is present. It is an interesting question for future work to determine the amount by which the system can be “pushed” off of the target orbit while preserving the transition.

### v2

@@ Acknowledgements.
We acknowledge helpful discussions with Wen-Wei Ho, Nishad Maskara, Lorenzo Piroli, and Sarang Gopalakrishnan. We especially thank Sriram Ganeshan and Michael Buchhold for numerous inspiring conversations and collaboration on related topics. This work was supported in part by the National Science Foundation under grants DMR-2143635 (T.I.) and DMR-2238895 (J.H.W.). J.H.P. was supported by the Office of Naval Research grant No. N00014-23-1-2357. This work was initiated and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation Grant No. PHY-1607611. This work was also supported in part by the International Centre for Theoretical Sciences (ICTS) via participation in the program “Periodically and quasi-periodically driven complex systems” (code: ICTS/pdcs2023/6). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (J.H.P.).
Appendix A Magnetization Measurement
Figure 14: (a) Performing the magnetization measurement, which involves a projector with a large rank can be thought of by controlling an ancillary system by the collective quantity of interest and then projectively measuring the ancillary system. (b) To update ancillary qubits, we conditionally apply $R_{j}$ to the $j$th ancilla. In the Fourier basis, this updates the ancillary state by $q_{i}$. (c) The full protocol to measure the magnetization of a system of $L=7$ qubits. It requires $3=\log_{2}(L+1)$ ancillas and an inverse Quantum Fourier Transform at the end to allow the measurement to be done in the $Z$-basis of each ancilla individually.
To measure magnetization in a system of qubits, we are inspired by Refs. [52, 54] (see also Refs. [53, 55]). The operator we are measuring is
$M=\sum_{i=1}^{L}Z_{i}.$
For ease in this section, we will use the equivalent number operator, where if $Z_{i}\ket{q_{i}}=(1-2q_{i})\ket{q_{i}}$ for $q_{i}=0,1$, then $n_{i}\ket{q_{i}}=q_{i}\ket{q_{i}}$. We want to accomplish the procedure of taking a state |$\ket{\Psi}$ and projecting it into the number sector $n$ using the projector
$P_{n}=\sum_{\sum_{i}q_{i}=n}\ket{q_{1}q_{2}\cdots q_{L}}\bra{q_{1}q_{2}\cdots q_{L}}$ |
with probability $\braket{\Psi}{P_{n}}{\Psi}$.
[…]
To accomplish this, we break it down into a few steps.
Step 1: Couple and measure an ancillary system. There are $L+1$ number states (00 to $L$), and thus, if we can couple in an $d>L$ dimensional qudit, we can projectively measure it to perform the collective measurement (illustrated in Fig. 14(a)).
To define this coupling, we need a unitary gate $U_{\mathrm{mag}}$ with the following property (where |$\ket{\cdot}_{a}$ represents the ancillary system)
$U_{\mathrm{mag}}\ket{q_{1}\cdots q_{L}}\otimes\ket{0}_{a}=\ket{q_{1}\cdots q_{L}}\otimes\big{\lvert}\sum_{i}q_{i}\big{\rangle}_{a}.$
[…]
with probability $\braket{\Psi}{P_{n}}{\Psi}$, achieving Eq. (18).
To accomplish this with quantum hardware, we can choose $d=2^{\lceil{\log_{2}(L+1)}\rceil}$, so that we have $\lceil\log_{2}(L+1)\rceil$ ancillary qubits.
Step 2: Designing $U_{\mathrm{mag}}$. There are a number of ways this can be designed, but here we describe one that uses the Fourier basis from the Quantum Fourier Transform [74].
Using $N=\lceil\log_{2}(L+1)\rceil$ qubits, define the Fourier states
|$\ket{Q}_{a}$ $\equiv\frac{1}{\sqrt{2^{N}}}\sum_{a_{1},\cdots a_{N}=0}^{1}e^{2\pi iQa/2^{N}}\ket{a_{1}}\otimes\cdots\otimes\ket{a_{N}},$ (20)
$=\frac{1}{\sqrt{2^{N}}}\bigotimes_{j=1}^{N}\left(\ket{0}+e^{2\pi iQ/2^{j}}\ket{1}\right).$ (21)
where $a_{j}=0,1$ and $a=a_{1}+a_{2}2^{1}+\cdots a_{N}2^{N-1}$ and thus $Q=0,\ldots,2^{N}-1$.
We therefore initialize the system in |$\ket{0}_{a}=\bigotimes_{j=1}^{N}\ket{+}$, and we can update the state using controlled phases between system qubit $i$ and ancilla $j$
$R_{m}(i,j)\ket{q_{i}}\otimes\ket{a_{j}}=e^{2\pi iq_{i}a_{j}/2^{m}}\ket{q_{i}}\otimes\ket{a_{j}},$ (22)
where all products should be understood to be right-to-left. By performing an $R_{j}(i,j)$ controlled phase on the $j$th ancilla, one can update the state
$\prod_{k=1}^{N}R_{k}(i,k)\ket{q_{i}}\otimes\ket{n}_{a}=\ket{q_{i}}\otimes\ket{n+q_{i}\!\!\!\mod d}_{a}.$
An example of this is in Fig. 14(b) labeled as the $+1$ gate (so named because it either leaves the ancilla register alone or adds one). If we subsequently apply this to all qubits, each gets added so that
$\prod_{i=1}^{L}\prod_{k=1}^{N}R_{k}(i,k)\ket{q_{1}\cdots q_{L}}\otimes\ket{0}_{a}=\ket{q_{1}\cdots q_{L}}\otimes\big{\lvert}\sum_{i}q_{i}\big{\rangle}_{a}.$
We have thus created exactly the unitary
$U_{\mathrm{mag}}=\prod_{i=1}^{L}\prod_{k=1}^{N}R_{k}(i,k).$
To measure the magnetization then, assuming we can only measure single qubits, we need to convert from the Fourier basis to the qubit basis, which is accomplished with an inverse Quantum Fourier Transform (IQFT) purely on the ancillas. If we then measure the ancillas, the readout will be the magnetization in binary. This full protocol for a system of size $L=7$ is illustrated in Fig. 14(c). (Note that similar circuits are presented in Refs. [53, 55], and that the approach outlined here is equivalent to quantum phase estimation for the operator $M$ applied to a state that is not an $M$ eigenstate.)
After this is done, the ancilla register can be re-used in either of two ways: Performing a Quantum Fourier transform again and tracking how it is updated by the next $U_{\mathrm{mag}}$ or resetting all ancillas into the $\ket{+}$ state to restart the procedure.
Appendix B Control onto General Short Orbits
@@
Appendix C Finite-Size Decay of Observables
@@
To remove this decay, we add one more step to the chaotic circuit of Sec. II.1. After applying $U_{\rm PXP}$ and $U_{\sigma}(\theta)$, we apply a (conditional) bit flip gate of the form (9) on a single qubit, chosen at random each time the chaotic circuit is applied; The rationale for this is that a finite fraction of the wave function gets trapped on the vacuum orbit and if we introduce a perturbation that is sub-extensive and kicks us off of the orbit, the system will not fully decay in the putative volume-law phase (at small system sizes). The result is illustrated with dashed lines in Fig. 16, which display clear saturation as expected in a standard dynamical phase. While we have not verified that the transition persists in the presence of this modification, a similar strategy was employed for $\beta$-adic Rényi circuits in Refs. [7, 8, 9, 10], where a clear second-order phase transition is present. It is an interesting question for future work to determine the amount by which the system can be “pushed” off of the target orbit while preserving the transition.